# Real-Time Doppler and Ionospheric Dispersion Correction Techniques for Arbitrary Waveforms Utilizing GPU Compute

### v1

@@ Appendix B Correction to Pseudo–Chebyshev Polynomial Expansion Method
The values of $T_{1}$ and $T_{2}$ shown in (22) and (23) only differ by using a factor of $4$ instead of $2$ with each $K_{2}$ term. We originally noticed this difference numerically by comparing parameters for the FFT ionospheric correction method presented in (15). We then repeated the original derivation from Ref. [1] as well as analytically derived an exact solution to (2), which we show in (9). We note that in Ref. [1] a substitution of a factor of $2$ is made for a factor of $4$, about which the authors state was determined “by simulation”. Without more information, we are unable to repeat their simulations to determine the validity of that substitution. However, when we rederived the Pseudo–Chebyshev method polynomial expansion, we arrived at the solutions shown in (7) and (8), which the FFT ionospheric correction method agrees with.
Figure 7: Frequency residuals of the Pseudo–Chebyshev method with our exact solution shown in (9) for a given LFM with $B=20~\text{MHz}$, $f_{c}=400~\text{MHz}$, and $T=0.1~\text{ms}$. Here, the corrected residual is the difference returned when using the coefficients in (7) and (8); the uncorrected curve using the coefficients in (22) and (23). Note that in this case the corrected Pseudo–Chebyshev method resulted in the lower error by over a full order of magnitude.
To further verify this change, we compared the original formulation of the Pseudo–Chebyshev method with our corrected coefficients, which is shown in Fig. 7. As shown in the plot, our corrected coefficient values applied for the given frequency resulted in a difference over a full order of magnitude lower than with the uncorrected method. This difference resulted in about $1~\text{dB}$ of SNR loss when cross correlated with our cubic solution. Based upon the errors from our exact analytic solution, comparison with our FFT-based ionospheric correction method, and analytic derivations, we conclude that the original factor of 2 difference presented in Ref. [1] was a slight error, which we have corrected here.
@@ Appendix C Algorithmic Optimizations for Sinc Interpolation on a GPU
For optimal performance, we elected to implement windowed Whittaker–Shannon interpolation on an NVIDIA H100 GPU connected to a $3.25~\text{GHz}$ CPU via PCIe 5.0. Our implementation was written in CUDA C++. Implementation on a GPU allowed us to parallelize a large portion of the computations. Here we will discuss three potential optimizations that were considered, their effect on performance in the H100, and considerations for use on different hardware.
Figure 8: Timing performance for this paper’s implementation of windowed Whittaker–Shannon interpolation times on an NVIDIA H100 GPU utilizing 32 blocks per kernel. Each plot demonstrates kernel completion times for four different implementations considered and ran for $10^{7}$ time trials. The input waveform contained $N=2^{19}~\text{samples}$ with a window size of $N_{window}=25~\text{samples}$. Figure 8a shows the interpolation time with no optimizations made outside of basic parallelization, which averaged completion around $150.5~\mu\text{s}$. Figure 8b displays the timing data if we first pre-load the input array into shared memory (tiling), which averaged completion around $159~\mu\text{s}$. Figures 8c and 8d present the timing performance when a LUT was included. The LUTs were size 3000 and 300 for the Figures 8c and 8d respectively. These kernels averaged completion in about $60.5~\mu s$ for the LUT size of 3000 and $59.3~\mu\text{s}$ for the LUT size of 300.
The primary optimization made during the implementation of this code was parallelization on a GPU. The standard method of parallelizing code on a GPU is to assign each executing thread to an index in the output array. This procedure is not only an efficient way to organize the workload on any given thread, but also assigning each thread a unique output memory address guarantees thread safety during the kernel call. In the case of windowed Whittaker–Shannon interpolation, each thread is responsible for the single summation of a window and writing the resultant interpolant to the output array. Recall from Table II that for an output array of $N$ elements, and a window size of $N_{window}$ elements, windowed sinc interpolation requires $\mathcal{O}\left(N\times N_{window}\right)$ time to complete. If the number of threads in the kernel meets or exceeds $N$, then the parallelized computation on a GPU reduces to $\mathcal{O}\left(N_{window}\right)$ time. For clarity, pseudo-code of this implementation can be seen in Algorithm 1 and we visually represented this parallelization in Fig. 9.
Algorithm 1 Parallelized Windowed Whittaker–Shannon Interpolation
@@
Figure 9: A visual representation of the data accesses during a summation of our parallelized implementation of windowed Whittaker–Shannon interpolation. In this parallelization, every thread in the GPU kernel is allocated to an output index $y_{n}$. For a given thread such that $n=d$, the output index will be $y_{d}$. That worker thread identifies the nearest samples to it in the original waveform, shown as $x_{c}$, which will be the center of the window. The thread then considered all values of $x_{n}$ in the window, which range from $x_{c-(N_{window}-1)/2}$ to $x_{c+(N_{window}-1)/2}$. After the summation, the total is written to the output index.
@@
Algorithm 2 Parallelized Windowed Whittaker–Shannon Interpolation w/ Lookup Table
@@
The final optimization we consider here is reduction in the number of read accesses to global memory from the input array of samples. Each thread in the kernel is responsible for $N_{window}$ global memory accesses—one for each element in the summation. Fortunately all threads within a block read from a highly-overlapping region of the input array. Therefore, global memory access can be reduced by allocating a region of shared memory for each block in the kernel, and then having each thread perform a single global memory read to initialize the input array into shared memory. This is a method of preloading the global memory into the block’s shared memory, referred to as “tiling”. Typically tiling is utilized to accelerate memory accesses in a 2D array (with each 2D region being a tile), but it can be applied in our 1D case. The major drawback of the use of tiling is the time cost of thread synchronization that must occur each time the shared memory is updated with a new region of input samples. We found that for our typical choice of interpolation parameters, tiling resulted in about a $9-10~\mu\text{s}$ slow-down in computational time. Because the number of memory accesses to the input array scales as $\mathcal{O}\left(N_{window}\right)$, tiling should become optimal for large enough choice of $N_{window}$, and may be optimal on another choice of hardware where global memory accesses are slower. Because tiling was slightly slower for our use case and for simplicity, we chose to exclude this potential optimization from our implementation of Whittaker–Shannon interpolation. Algorithm 3 presents pseudo-code for our version of sinc interpolation with tiling.
Algorithm 3 Parallelized Windowed Whittaker–Shannon Interpolation w/ Tiling
@@
Note that the implementations of a LUT and tiling both require shared memory allocation that is related to $N_{window}$ in size. It is possible that tiling and the use of a LUT will be incompatible on devices with sufficiently limited shared memory and/or a sufficiently large choice of $N_{window}$. Based on our timing performance and because the number of sinc value calculations scales as $\mathcal{O}\left(N_{window}\right)$, it seems likely that the implementation of a LUT will most-likely be optimal in any situation where a LUT and tiling are mutually exclusive.
@@
For completeness, we note that a great deal of work has been done on efficient Whittaker–Shannon interpolation by padding elements in the frequency domain via the use of DFTs [24, 25]. For efficient processing, these methods require that the number of interpolants be proportional to the number of input samples by a power of 2. This condition is compatible with applications where the number of interpolants, $N_{I}$, is selectable such as image processing. When this condition is met, Whittaker–Shannon interpolation is achievable with time complexities of $\mathcal{O}\left(N_{I}\log{N_{I}}\right)$. Interpolation for Doppler dispersion correction requires a specific choice of $N_{I}$ that depends upon $N$ and the target velocity. The variable number of interpolants results in variable time complexity to complete the required FFTs. For our use case, these methods resulted in an unpredictable signal processing rate that was consistently slower than all implementations shown in Fig. 8—in some cases by a factor of $10\times$ slower rates. These trade offs are extremely similar to considerations for the FFT P/Q resampling method. We conclude that for radar signal processing, performing the interpolation in the time domain directly from (16) is often more efficient and always more predictable in terms of kernel completion times.

### v2

@@ Appendix B Correction to Pseudo-Chebyshev Polynomial Expansion Method
The values of $T_{1}$ and $T_{2}$ shown in (23) and (24) only differ by using a factor of $4$ instead of $2$ with each $K_{2}$ term. We originally noticed this difference numerically by comparing parameters for the FFT ionospheric correction method presented in (16). We then repeated the original derivation from Ref. [1] as well as analytically derived an exact solution to (4), which we show in (11). We note that in Ref. [1] a substitution of a factor of $2$ is made for a factor of $4$, about which the authors state was determined “by simulation”. Without more information, we are unable to repeat their simulations to determine the validity of that substitution. However, when we rederived the Pseudo-Chebyshev method polynomial expansion, we arrived at the solutions shown in (9) and (10), which the FFT ionospheric correction method agrees with.
Figure 8: Frequency residuals of the Pseudo-Chebyshev method with our exact solution shown in (11) for a given LFM with $B=20~\text{MHz}$, $f_{c}=400~\text{MHz}$, and $T=0.1~\text{ms}$. Here, the corrected residual is the difference returned when using the coefficients in (9) and (10); the uncorrected curve using the coefficients in (23) and (24). Note that in this case the corrected Pseudo-Chebyshev method resulted in the lower error by over a full order of magnitude.
To further verify this change, we compared the original formulation of the Pseudo-Chebyshev method with our corrected coefficients, which is shown in Fig. 8. As shown in the plot, our corrected coefficient values applied for the given frequency resulted in a difference over a full order of magnitude lower than with the uncorrected method. This difference resulted in about $1~\text{dB}$ of SNR loss when cross correlated with our cubic solution. Based upon the errors from our exact analytic solution, comparison with our FFT-based ionospheric correction method, and analytic derivations, we conclude that the original factor of 2 difference presented in Ref. [1] was a slight error, which we have corrected here.
@@ Appendix C Algorithmic Optimizations for Sinc Interpolation on a GPU
For optimal performance, we elected to implement windowed Whittaker-Shannon interpolation on an NVIDIA H100 GPU connected to a $3.25~\text{GHz}$ CPU via PCIe 5.0. Our implementation was written in CUDA C++. Implementation on a GPU allowed us to parallelize a large portion of the computations. Here we will discuss three potential optimizations that were considered, their effect on performance in the H100, and considerations for use on different hardware.
((a))
((b))
((c))
((d))
Figure 9: Benchmarks for this paper’s implementation of windowed Whittaker-Shannon interpolation times on an NVIDIA H100 GPU utilizing 32 blocks per kernel. Each plot demonstrates kernel completion times for four different implementations considered and ran for $10^{7}$ time trials. The input waveform contained $N=2^{19}~\text{samples}$ with a window size of $N_{window}=25~\text{samples}$. Figure 9(a) shows the interpolation time with no optimizations made outside of basic parallelization, which averaged completion around $150.5~\mu\text{s}$. Figure 9(b) displays the timing data if we first pre-load the input array into shared memory (tiling), which averaged completion around $159~\mu\text{s}$. Figures 9(c) and 9(d) present the benchmarking when a LUT was included. The LUTs were size 3000 and 300 for the Figures 9(c) and 9(d) respectively. These kernels averaged completion in about $60.5~\mu s$ for the LUT size of 3000 and $59.3~\mu\text{s}$ for the LUT size of 300.
The primary optimization made during the implementation of this code was parallelization on a GPU. The standard method of parallelizing code on a GPU is to assign each executing thread to an index in the output array. This procedure is not only an efficient way to organize the workload on any given thread, but also assigning each thread a unique output memory address guarantees thread safety during the kernel call. In the case of windowed Whittaker-Shannon interpolation, each thread is responsible for the single summation of a window and writing the resultant interpolant to the output array. Recall from Table I that for an output array of $N$ elements, and a window size of $N_{window}$ elements, windowed sinc interpolation requires $\mathcal{O}\left(N\times N_{window}\right)$ time to complete. If the number of threads in the kernel meets or exceeds $N$, then the parallelized computation on a GPU reduces to $\mathcal{O}\left(N_{window}\right)$ time. For clarity, pseudo-code of this implementation can be seen in Algorithm 1 and we visually represented this parallelization in Fig. 10.
Algorithm 1 Parallelized Windowed Whittaker-Shannon Interpolation
@@
Figure 10: A visual representation of the data accesses during a summation of our parallelized implementation of windowed Whittaker-Shannon interpolation. In this parallelization, every thread in the GPU kernel is allocated to an output index $y_{n}$. For a given thread such that $n=d$, the output index will be $y_{d}$. That worker thread identifies the nearest samples to it in the original waveform, shown as $x_{c}$, which will be the center of the window. The thread then considered all values of $x_{n}$ in the window, which range from $x_{c-(N_{window}-1)/2}$ to $x_{c+(N_{window}-1)/2}$. After the summation, the total is written to the output index.
@@
Algorithm 2 Parallelized Windowed Whittaker-Shannon Interpolation w/ Lookup Table
@@
The final optimization we consider here is reduction in the number of read accesses to global memory from the input array of samples. Each thread in the kernel is responsible for $N_{window}$ global memory accesses—one for each element in the summation. Fortunately all threads within a block read from a highly-overlapping region of the input array. Therefore, global memory access can be reduced by allocating a region of shared memory for each block in the kernel, and then having each thread perform a single global memory read to initialize the input array into shared memory. This is a method of preloading the global memory into the block’s shared memory, referred to as “tiling”. Typically tiling is utilized to accelerate memory accesses in a 2D array (with each 2D region being a tile), but it can be applied in our 1D case. The major drawback of the use of tiling is the time cost of thread synchronization that must occur each time the shared memory is updated with a new region of input samples. We found that for our typical choice of interpolation parameters, tiling resulted in about a $9-10~\mu\text{s}$ slow-down in computational time. Because the number of memory accesses to the input array scales as $\mathcal{O}\left(N_{window}\right)$, tiling should become optimal for large enough choice of $N_{window}$, and may be optimal on another choice of hardware where global memory accesses are slower. Because tiling was slightly slower for our use case and for simplicity, we chose to exclude this potential optimization from our implementation of Whittaker-Shannon interpolation. Algorithm 3 presents pseudo-code for our version of sinc interpolation with tiling.
Algorithm 3 Parallelized Windowed Whittaker-Shannon Interpolation w/ Tiling
@@
Note that the implementations of a LUT and tiling both require shared memory allocation that is related to $N_{window}$ in size. It is possible that tiling and the use of a LUT will be incompatible on devices with sufficiently limited shared memory and/or a sufficiently large choice of $N_{window}$. Based on our benchmarks and because the number of sinc value calculations scales as $\mathcal{O}\left(N_{window}\right)$, it seems likely that the implementation of a LUT will most-likely be optimal in any situation where a LUT and tiling are mutually exclusive.
@@
For completeness, we note that a great deal of work has been done on efficient Whittaker-Shannon interpolation by padding elements in the frequency domain via the use of DFTs [17, 18]. For efficient processing, these methods require that the number of interpolants be proportional to the number of input samples by a power of 2. This condition is compatible with applications where the number of interpolants, $N_{I}$, is selectable such as image processing. When this condition is met, Whittaker-Shannon interpolation is achievable with time complexities of $\mathcal{O}\left(N_{I}\log{N_{I}}\right)$. Interpolation for Doppler dispersion correction requires a specific choice of $N_{I}$ that depends upon $N$ and the target velocity. The variable number of interpolants results in variable time complexity to complete the required FFTs. For our use case, these methods resulted in an unpredictable signal processing rate that was consistently slower than all implementations shown in Fig. 9—in some cases by a factor of $10\times$ slower rates. These trade offs are extremely similar to considerations for the FFT P/Q resampling method. We conclude that for radar signal processing, performing the interpolation in the time domain directly from (17) is often more efficient and always more predictable in terms of kernel completion times.